# Study of the decay pattern of $f_{0}(1370)$ as a $\kappa\bar{\kappa}$ molecular state

### v1

@@ II Formalism
In this section, we give the detailed formalism for calculating the partial decay widths of $f_{0}(1370)$ as a $\kappa\bar{\kappa}$ molecular state. Using the isospin conventions shown in Appendix A, we have
@@
Note that $\kappa$ is a well-known broad resonance and considered as a $K\pi$ molecule in some models, hence, considering two broad states to form a molecular state is counter-intuitive. However, this system can be essentially regarded as a $(K\pi)(\bar{K}\pi)$ four-body interacting system. To avoid the complexity of four-body dynamics, we argue that describing the system as a $\kappa\bar{\kappa}$ molecule is a reasonable and practical approximation. Under such a scenario, we attempt to test whether the decay pattern of the $f_{0}(1370)$ is consistent with the current data.
Firstly, we attempt to use the Weinberg compositeness to estimate the coupling of $f_{0}(1370)$ decay to the $\kappa\bar{\kappa}$ pair. The compositeness concept in terms of the hadronic degrees of freedom is model-independently related to the experimental observables for the S-wave interaction in the weak-binding limit, and can be generalized to the resonances [53, 54, 55]. In particular, it was shown that the field renormalization constant $Z$ of a weakly bound state can be related to the threshold parameters (the scattering length and the effective range) in a model-independent way [56].
The field renormalization constant $Z$ represents the probability of finding the elementary component $|B_{0}\rangle$ in the physical state $|B\rangle$, and the quantity X=1ZX=1-Z stands for the compositeness of the state. The criterion to determine the effective coupling based on binding energy and compositeness parameter $X$ is expressed as
@@
Figure 1: The mechanism for $f_{0}(1370)$ decay as a $\kappa\bar{\kappa}$ molecule state. (a) $\pi^{+}\pi^{-}$$\eta\eta$ final states. (b) $K^{+}K^{-}$ final states. (c) $4\pi$ final states. $S/V$ indicates the $t$-exchange meson could be $\kappa$ or $K^{*}$. (d) $K\bar{K}\pi\pi$ final states.
In the hadronic molecule picture, the schematic diagrams of $f_{0}(1370)$ decay to different channels, including $\pi\pi$, $K\bar{K}$, $\eta\eta$, $4\pi$ and $K\bar{K}\pi\pi$, are shown in Fig. 1. Note that the vertices with quantum numbers $SSP$, SPVSPV are forbidden and with quantum numbers $SSS$, SSVSSV, SPPSPP, SVVSVV are allowed due to the parity conservation laws. The effective Lagrangians adopted for these allowed vertices have the forms as
@@
Notice that there are several distinct mechanisms contributing to the same final states, especially for the $4\pi$ channel, hence we should explicitly care about the phase of the couplings in order to pin down the interference pattern. Expanding the Lagrangian in Eq. (6) and focusing on the $\sigma\pi\pi$ vertex, we can obtain:
$\mathcal{L}_{\sigma\pi\pi}=\frac{g_{SPP}}{\sqrt{2}}\cdot\sigma\pi^{0}\pi^{0}+\sqrt{2}g_{SPP}\cdot\sigma\pi^{+}\pi^{-},$ (10)
@@
note that due to the identical feature, there are two ways to contract the two $\pi^{0}$ fields and the two final $\pi^{0}$ states for the first term. Based on this effective Lagrangian, the effective couplings are defined as follows: $g_{\sigma\pi^{0}\pi^{0}}=g_{SPP}/\sqrt{2}$ and $g_{\sigma\pi^{+}\pi^{-}}=\sqrt{2}g_{SPP}$. The effective couplings $g_{\kappa K\pi}$ and $g_{\sigma\pi\pi}$ we define here can be related to the residues at the resonance poles of $\kappa$ in $K\pi$ scattering and $\sigma$ in $\pi\pi$ scattering, respectively. We adopt the residue $\tilde{g}_{\sigma\pi\pi}=3.61e^{-1.30i}$ GeV for the $\sigma$ pole from Ref. [59, 60]. By transforming this coupling for $\pi\pi$ with fixed isospin $I=0$ to the coupling associated with specific particles based on Eq. (43), we have $g_{\sigma\pi^{+}\pi^{-}}=\tilde{g}_{\sigma\pi\pi}(-\sqrt{2}/\sqrt{3})=-0.79+2.84i$ GeV and $g_{\sigma\pi^{0}\pi^{0}}=g_{\sigma\pi^{+}\pi^{-}}/2=-0.39+1.42i$ GeV. Ref. [59] also provides the coupling $\tilde{g}_{\rho\pi\pi}=6.01e^{-0.09i}$.
@@
We adopt the residues $\tilde{g}_{\kappa^{0}K\pi}=3.81e^{-1.49i}$ GeV and $\tilde{g}_{K^{*0}K\pi}=5.69e^{-0.076i}$ from Ref. [61]. Similarly, they can be transferred to the couplings associated with specific particles based on Eq. (35): $g_{\kappa^{0}K^{+}\pi^{-}}=\tilde{g}_{\kappa^{0}K\pi}(\sqrt{2}/{\sqrt{3}})=0.25-3.10i$ and $g_{K^{*0}K^{+}\pi^{-}}=\tilde{g}_{K^{*0}K^{+}\pi^{-}}(\sqrt{2}/{\sqrt{3}})=4.63-0.35i$. One can find that due to the pole positions of $\rho$ and $K^{*}$ are relatively close to the real axis, their corresponding residues have small argument and are almost real numbers. The strength of coupling $g_{K^{*0}K^{+}\pi^{-}}$ ($4.64$) is very closed to the strength of $g_{\phi K^{+}K^{-}}$ extracted from the decay of $\phi\to K\bar{K}$, e.g., $|g_{\phi K^{+}K^{-}}|=4.47$ is determined in Ref. [62]. Such phenomenon is a confirmation of the flavor symmetry result that $g_{K^{*0}K^{+}\pi^{-}}=g_{\phi K^{+}K^{-}}=g_{VPP}$.
@@
Now come to discuss the remaining relevant couplings. The module of $g_{\sigma\pi^{+}\pi^{-}}$ adopted here is $2.95$. Similarly, from the expanded Lagrangian we adopt that $g_{\sigma K\bar{K}}=g_{\sigma\pi^{+}\pi^{-}}/2$ under the SU(3) symmetry, leading to the ratio of $g_{\sigma K\bar{K}}/g_{\sigma\pi^{+}\pi^{-}}$ is $0.5$. These couplings are also extracted in Ref. [63] by fitting the $\pi\pi\to\pi\pi/K\bar{K}$ scattering data with different models of $K$- and $S$-matrices. They found the average value of $|g_{\sigma\pi^{+}\pi^{-}}|=2.47\pm 0.45$ GeV and about $0.6\pm 0.2$ for this ratio, both quantities are consistent with the our adoptive values.
@@
The $\sigma\kappa\kappa$ coupling can be approximately related to the $\sigma K\bar{K}$ and $\sigma\pi\pi$ couplings when we treat this vertex as a decay of $\kappa\to\kappa\sigma$ where the component $\pi$ or $K$ of $\kappa$ can be considered as a spectator. Analogous estimations can be applied to the coupling $g_{\sigma\kappa\bar{K}^{*}}$, which can be related to the couplings $g_{K^{*0}K^{+}\pi^{-}}$, and the coupling $\rho\kappa\bar{\kappa}$ which can be related to the couplings $\rho\pi\pi$ and $\rho K\bar{K}$. The $\rho K^{*}\bar{\kappa}$ coupling can be related to another $VVS$ type coupling $\sigma\rho\rho$ under flavor symmetry with coupling $g_{\sigma\rho^{+}\rho^{-}}$ determined by assuming $g_{\sigma\pi^{+}\pi^{-}}=g_{\sigma\rho^{+}\rho^{-}}$. For convenience, we collect the explicit formulae for estimating these couplings in Appendix B and the accurate values of these couplings in Table 1.
@@
Based on these effective Lagrangians, the amplitude of the diagram in Fig 1(a) decaying to $\pi^{+}\pi^{-}$ via the neutral $\kappa^{0}\bar{\kappa}^{0}$ component is written as
@@
Similarly, we can obtain the amplitude in this mechanism via the charged $\kappa^{+}\kappa^{-}$ rescattering and denoted it as $\mathcal{M}^{C}_{\pi^{+}\pi^{-}}$. Note that in the $\pi^{+}\pi^{-}$ channel no charge conjugate loop can contribute for both the neutral and charged $\kappa$ components. The kinematic variables of the particles in the triangle loop are also denoted in Fig. 1(a). The effective Lagrangian approach, which deals with the nonlocal effects by introducing empirical form factors to cut off the divergences, are applied to broader kinematic regions. Hence, we include a commonly adopted monopole form factor to the exchanged mesons [64],
@@
where $\Lambda=m_{\text{ex}}+\alpha\Lambda_{\text{QCD}}$ with $\Lambda_{\text{QCD}}=300~\text{MeV}$ and $\alpha=1\sim 2$ is widely chosen as a cut-off parameter. $p_{\text{ex}}$ and $m_{\text{ex}}$ are the four-momentum and mass of the exchanged meson, respectively. However, the disadvantage is that, due to lack of knowledge about the behavior of the counter terms, model dependence will be present in association with the cutoff energies and different forms for the form factors [65].
@@
where $\Phi_{\pi\pi}$ indicate the $\pi\pi$ phase space. The factor $3/2$ comes from the inclusion of the $\pi^{0}\pi^{0}$ final state considering their identical feature and the relations in couplings.
The mechanism of $f_{0}(1370)$ decay to $\eta\eta$ is the same as the decay to $\pi^{+}\pi^{-}$, where just changing the vertices of $\kappa\to K\pi$ to $\kappa\to K\eta$. Hence, by replacing these couplings we can obtain the amplitudes $\mathcal{M}^{C}_{\eta\eta}$ and $\mathcal{M}^{N}_{\eta\eta}$ for the $\eta\eta$ final states. Note that the charge conjugate loop can contribute for both the neutral and charged $\kappa$ components in this channel. Then we obtain the decay width of the $\eta\eta$ channel by
@@
$\Phi_{\eta\eta}$ is the phase space of $\eta\eta$ final states and $1/2$ is the identical factor.
@@
The coupling $g_{\kappa K\eta}$ can be related to $g_{\kappa^{0}K^{+}\pi^{-}}$ by the flavor symmetry. In fact the physical $\eta$ state is the mixing of $\eta_{8}$ and $\eta_{1}$ with $\eta=\cos\theta\eta_{8}-\sin\theta\eta_{1}$ due to the SU(3) flavor breaking. As shown in Table 1, we adopt two strategies to deal with the $g_{\kappa K\eta}$ coupling: one is that we ignore the $\eta_{1}$ component in $\eta$ by taking $\eta=\cos\theta\eta_{8}$, and restrict to the symmetry inside the octet by using $P_{8}$ as the pseudoscalar matrix. In this case, we denote this coupling as $g_{\kappa K\eta}$; the other case is that we take $\eta=\cos\theta\eta_{8}-\sin\theta\eta_{1}$ and relate the couplings associated with these two part as well as the $g_{\kappa K\pi}$ by the U(3) flavor symmetry, where the pseudoscalar matrix is taken as $U$. In this case, this coupling is denoted as $g^{\prime}_{\kappa K\eta}$. Here we should be aware of that the U(3) symmetry is not so good in the pseudoscalar sector due to the U(1) anomaly. The explicit forms of $P_{8}$ and $U$ can be found in Appendix. C .
@@
Same as the case in $\pi\pi$ channel, there is no charge conjugate loop that can contribute. Considering the relation between the couplings $g_{{\kappa^{0}}K^{+}\pi^{-}}$ and $g_{{\kappa^{+}}K^{+}\pi^{0}}$, we found that $\mathcal{M}^{C}_{K^{+}K^{-}}=(1/2)\mathcal{M}^{N}_{K^{+}K^{-}}$. The amplitude of $f_{0}(1370)$ decay to $K^{+}K^{-}$ channel by exchanging $\eta$ is written as
@@
where the $\Phi_{K\bar{K}}$ indicate the $K\bar{K}$ phase space. The factor $2$ comes from the inclusion of $K^{0}\bar{K}^{0}$ final state.
For the $4\pi$ decay channel, there are three different specific final states which are $2\pi^{+}2\pi^{-}$, $\pi^{+}\pi^{-}2\pi^{0}$ and $4\pi^{0}$ respectively. As claimed by the analyses in experiments [20, 23, 24], the $4\pi$ final states are mainly from the intermediate channels of $\sigma\sigma$ (i.e. $f_{0}(500)$) and $\rho\rho$. This mechanism is shown in Fig. 1(c). Specifically, for the initial decay of an isoscalar state $f_{0}$, the $2\pi^{+}2\pi^{-}$ final state can only come from the $\rho^{0}\rho^{0}$ and $\sigma\sigma$ intermediate channels; the $\pi^{+}\pi^{-}2\pi^{0}$ final state can only come from the $\rho^{+}\rho^{-}$ and $\sigma\sigma$ intermediate channels; and the $4\pi^{0}$ final state can only come from the $\sigma\sigma$ intermediate channel. Note that, due to the identical feature, the exchange of the identical particles can also produce different diagrams that can contribute.
Firstly we focus on the decay width of the $2\pi^{+}2\pi^{-}$ channel, then the decay width of the other two channels can be related to $\Gamma_{2\pi^{+}2\pi^{-}}$ by considering the relations among the couplings in the isospin limit. As shown in Fig. 1(c), we consider the decay mechanism of $\kappa\bar{\kappa}$ rescattering to $\rho\rho$ or $\sigma\sigma$ first and subsequently decaying to $4\pi$. The amplitude of the process for $\kappa\bar{\kappa}$ rescattering to $2\sigma\to 2\pi^{+}2\pi^{-}$ by exchanging scalar $\kappa$, denoted as $\mathcal{M}^{S}_{\sigma}$, has the form of:
@@
momenta of the final $\pi$ are denoted as $\pi^{+}({p_{a}})\pi^{-}(p_{b})\pi^{+}(p_{c})\pi^{-}(p_{d})$, and $p_{ab}=p_{a}+p_{b}$, $p_{cd}=p_{c}+p_{d}$ are respectively the momenta of the two intermediate $\sigma$. The $ex.$ term indicates the contribution from exchanging the identical particles, which is the same as the first term but only change the momenta of the two $\sigma$ to $p_{ad}=p_{a}+p_{d}$ and $p_{bc}=p_{b}+p_{c}$ respectively. Note that, for such a decay mechanism, there exist two charge conjugate loops for both the charged and neutral $\kappa$ components. By analyzing the relations of the couplings involved, we find that the amplitudes of these four diagrams are identical, resulting in a multiplicative factor of $4$.
@@
The amplitude of the process for the $\kappa^{0}\bar{\kappa}^{0}$ component rescattering to $2\rho^{0}\to 2\pi^{+}2\pi^{-}$ by exchanging $\kappa^{0}$, as shown in Fig. 1(c), has the form of
@@
The amplitude of the process for the $\kappa^{0}\bar{\kappa}^{0}$ component rescattering to $2\rho^{0}\to 2\pi^{+}2\pi^{-}$ by exchanging $K^{*0}$, as shown in Fig. 1(c), has the form of
@@
here $G=g_{F\kappa\bar{\kappa}^{0}}\cdot g^{2}_{\rho^{0}\kappa^{0}\bar{K}^{*0}}\cdot g^{2}_{\rho^{0}\pi^{+}\pi^{-}}$. The total amplitude of decaying to $2\pi^{+}2\pi^{-}$ is the sum of all the amplitudes above, i.e.
@@
The constant $1/4$ are the symmetry factor on account of the two kinds of identical particles: two $\pi^{+}$ and two $\pi^{-}$.
Then, we could discuss the decay amplitudes of $\pi^{+}\pi^{-}2\pi^{0}$ and $4\pi^{0}$ states based on the amplitude of the $2\pi^{+}2\pi^{-}$ state. For the $\rho^{+}\rho^{-}\to\pi^{+}(p_{a})\pi^{0}(p_{b})$-$\pi^{-}(p_{c})\pi^{0}(p_{d})$ process, we find the following similarities and differences compared to the $\rho^{0}\rho^{0}\to\pi^{+}(p_{a})\pi^{-}(p_{b})$-$\pi^{+}(p_{c})\pi^{-}(p_{d})$:
  • There is only one kind of loop for both the charged and neutral $\kappa$ components.
@@
Considering the above items we found that the amplitudes adapted to the $\rho^{+}\rho^{-}\to\pi^{+}\pi^{-}2\pi^{0}$ process are just $\mathcal{M}^{S}_{\rho}$ and $\mathcal{M}^{V}_{\rho}$.
For the $\sigma\sigma\to\pi^{+}(p_{a})\pi^{-}(p_{c})$-$\pi^{0}(p_{b})\pi^{0}(p_{d})$ process, because the two identical $\pi^{0}$ come from the same $\sigma$, the exchange between them are topologically identical. Note that the relation $g_{\sigma\pi^{+}\pi^{-}}=2g_{\sigma\pi^{0}\pi^{0}}$, then, the amplitudes $\mathcal{M}^{S}_{\sigma}$ and $\mathcal{M}^{V}_{\sigma}$ adapted to the $\pi^{+}\pi^{-}2\pi^{0}$ state are obtained by changing the momentums of the two $\sigma$ to $p_{ac}=p_{a}+p_{c}$ and $p_{bd}=p_{b}+p_{d}$ respectively and eliminating the exchange terms.

### v2

@@ II Formalism
In this section, we present the detailed formalism for calculating the partial decay widths of the $f_{0}(1370)$ as a $\kappa\bar{\kappa}$ molecular state. Using the isospin conventions shown in Appendix A, we have
@@
Given that the $\kappa$ is a broad resonance often modeled as a $K\pi$ molecule, considering two such broad states to form a molecular state is counter-intuitive. However, this system can be essentially regarded as a $(K\pi)(\bar{K}\pi)$ four-body interacting system. To avoid the complexity of four-body dynamics, we argue that describing the system as a $\kappa\bar{\kappa}$ molecule is a reasonable and practical approximation. Within this framework, we evaluate whether the predicted decay pattern of the $f_{0}(1370)$ remains consistent with available experimental data.
Firstly, we attempt to use the Weinberg compositeness to estimate the coupling of $f_{0}(1370)$ decay to the $\kappa\bar{\kappa}$ pair. The compositeness concept in terms of the hadronic degrees of freedom is model-independently related to the experimental observables for the $S$-wave interaction in the weak-binding limit, and can be generalized to the resonances [53, 54, 55]. In particular, the field renormalization constant $Z$ of a weakly bound state can be related to the threshold parameters (the scattering length and the effective range) in a model-independent way  [56]. The $Z$ represents the probability of finding the elementary component $|B_{0}\rangle$ in the physical state $|B\rangle$, and the quantity X=1ZX=1-Z stands for the compositeness of the state. The criterion to determine the effective coupling based on binding energy and compositeness parameter $X$ is expressed as
@@
Figure 1: The mechanism for $f_{0}(1370)$ decay as a $\kappa\bar{\kappa}$ molecular state. (a) $\pi^{+}\pi^{-}$$\eta\eta$ final states. (b) $K^{+}K^{-}$ final states. (c) $4\pi$ final states. $S/V$ denotes that the $t$-channel exchange meson is either a scalar $\kappa$ or a vector $K^{*}$. (d) $K\bar{K}\pi\pi$ final states.
In the hadronic molecule picture, the schematic diagrams of $f_{0}(1370)$ decay to different channels, including $\pi\pi$, $K\bar{K}$, $\eta\eta$, $4\pi$ and $K\bar{K}\pi\pi$, are shown in Fig. 1. Note that the vertices with quantum numbers $SSP$, SPVSPV are forbidden and with quantum numbers $SSS$, SSVSSV, SPPSPP, SVVSVV are allowed based on the parity conservation laws. For the allowed vertices, the adopted effective Lagrangians have the forms as
@@
It is important to note that several distinct mechanisms contribute to the same final states, particularly in the $4\pi$ channel. Consequently, the relative phases of the couplings must be explicitly taken into account to determine the interference pattern. Expanding the Lagrangian in Eq. (6) and focusing on the $\sigma\pi\pi$ vertex, we can obtain:
$\mathcal{L}_{\sigma\pi\pi}=\frac{g_{SPP}}{\sqrt{2}}\cdot\sigma\pi^{0}\pi^{0}+\sqrt{2}g_{SPP}\cdot\sigma\pi^{+}\pi^{-}.$ (10)
@@
Note that due to the identical feature, there are two ways to contract the two $\pi^{0}$ fields and the two final $\pi^{0}$ states for the first term. Based on this effective Lagrangian, the effective couplings are defined as follows: $g_{\sigma\pi^{0}\pi^{0}}=g_{SPP}/\sqrt{2}$ and $g_{\sigma\pi^{+}\pi^{-}}=\sqrt{2}g_{SPP}$. The effective couplings $g_{\kappa K\pi}$ and $g_{\sigma\pi\pi}$ we define here can be related to the residues at the resonance poles of $\kappa$ in $K\pi$ scattering and $\sigma$ in $\pi\pi$ scattering, respectively. We adopt the residue $\tilde{g}_{\sigma\pi\pi}=3.61e^{-1.30i}$ GeV for the $\sigma$ pole from Refs. [59, 60]. By transforming this coupling for $\pi\pi$ with fixed isospin $I=0$ to the coupling associated with specific particles based on Eq. (43), we have $g_{\sigma\pi^{+}\pi^{-}}=\tilde{g}_{\sigma\pi\pi}(-\sqrt{2}/\sqrt{3})=-0.79+2.84i$ GeV and $g_{\sigma\pi^{0}\pi^{0}}=g_{\sigma\pi^{+}\pi^{-}}/2=-0.39+1.42i$ GeV. Ref. [59] also provides the coupling $\tilde{g}_{\rho\pi\pi}=6.01e^{-0.09i}$.
@@
We adopt the residues $\tilde{g}_{\kappa^{0}K\pi}=3.81e^{-1.49i}$ GeV and $\tilde{g}_{K^{*0}K\pi}=5.69e^{-0.076i}$ from Ref. [61]. Similarly, they can be transferred to the couplings associated with specific particles based on Eq. (35): $g_{\kappa^{0}K^{+}\pi^{-}}=\tilde{g}_{\kappa^{0}K\pi}(\sqrt{2}/{\sqrt{3}})=0.25-3.10i$ and $g_{K^{*0}K^{+}\pi^{-}}=\tilde{g}_{K^{*0}K^{+}\pi^{-}}(\sqrt{2}/{\sqrt{3}})=4.63-0.35i$. Because the pole positions of $\rho$ and $K^{*}$ are relatively close to the real axis, their corresponding residues have small argument and are almost real numbers. The strength of coupling $|g_{K^{*0}K^{+}\pi^{-}}|=4.64$ is very close to the strength of $g_{\phi K^{+}K^{-}}$ extracted from the decay of $\phi\to K\bar{K}$, e.g., $|g_{\phi K^{+}K^{-}}|=4.47$ determined in Ref. [62]. Such a phenomenon is a confirmation of the flavor symmetry result that $g_{K^{*0}K^{+}\pi^{-}}=g_{\phi K^{+}K^{-}}=g_{VPP}$.
@@
Now we turn to the discussion of the remaining relevant couplings. The magnitude of $g_{\sigma\pi^{+}\pi^{-}}$ adopted here is $2.95$. Similarly, from the expanded Lagrangian we adopt that $g_{\sigma K\bar{K}}=g_{\sigma\pi^{+}\pi^{-}}/2$ under the SU(3) symmetry, leading to the ratio of $g_{\sigma K\bar{K}}/g_{\sigma\pi^{+}\pi^{-}}$ is $0.5$. These couplings are also extracted in Ref. [63] by fitting the $\pi\pi\to\pi\pi/K\bar{K}$ scattering data with different models of $K$- and $S$-matrices. They found the average value of $|g_{\sigma\pi^{+}\pi^{-}}|=2.47\pm 0.45$ GeV and about $0.6\pm 0.2$ for the ratio of $g_{\sigma K\bar{K}}/g_{\sigma\pi^{+}\pi^{-}}$, both quantities are consistent with the our adopted values.
@@
The $\sigma\kappa\kappa$ coupling can be approximately related to the $\sigma K\bar{K}$ and $\sigma\pi\pi$ couplings when the component $\pi$ or $K$ of the $\kappa$ is considered as a spectator. Analogous estimations can be applied to the coupling $g_{\sigma\kappa\bar{K}^{*}}$ which can be related to the couplings $g_{K^{*0}K^{+}\pi^{-}}$, and the coupling $\rho\kappa\bar{\kappa}$ which can be related to the couplings $\rho\pi\pi$ and $\rho K\bar{K}$. The $\rho K^{*}\bar{\kappa}$ coupling can be related to another $VVS$ type coupling $\sigma\rho\rho$ under flavor symmetry with coupling $g_{\sigma\rho^{+}\rho^{-}}$ determined by assuming $g_{\sigma\pi^{+}\pi^{-}}=g_{\sigma\rho^{+}\rho^{-}}$. For convenience, we collect the explicit formulae for estimating these couplings in Appendix B and list their specific values in Table 1.
@@
Based on these effective Lagrangians, the amplitude of the diagram in Fig. 1(a) decaying to $\pi^{+}\pi^{-}$ via the neutral $\kappa^{0}\bar{\kappa}^{0}$ component is written as
@@
Similarly, we can obtain the amplitude in this mechanism via the charged $\kappa^{+}\kappa^{-}$ component and denoted it as $\mathcal{M}^{C}_{\pi^{+}\pi^{-}}$. Notice that in the $\pi^{+}\pi^{-}$ channel no charge conjugated loop can contribute for both the neutral and charged $\kappa$ components. The kinematic variables of the particles in the triangle loop are also denoted in Fig. 1(a). The effective Lagrangian approach, which deals with the nonlocal effects by introducing empirical form factors to cut off the divergences, are applied to broader kinematic regions. Hence, we include a commonly adopted monopole form factor to the exchanged mesons [64],
@@
where $\Lambda=m_{\text{ex}}+\alpha\Lambda_{\text{QCD}}$ with $\Lambda_{\text{QCD}}=300~\text{MeV}$ and $\alpha=1\sim 2$ is widely chosen as a cut-off parameter. $p_{\text{ex}}$ and $m_{\text{ex}}$ are the four-momentum and mass of the exchanged meson, respectively. However, due to the lack of knowledge about the behavior of the counter terms, model dependence will be present in association with the cutoff energies and different forms for the form factors [65].
@@
where $\Phi_{\pi\pi}$ indicate the $\pi\pi$ phase space. The factor $3/2$ originates from the inclusion of the $\pi^{0}\pi^{0}$ final state considering their identical feature and the relations in couplings.
The mechanism for the $f_{0}(1370)$ decay to $\eta\eta$ is the same as that for decay to $\pi^{+}\pi^{-}$, where just changing the vertices of $\kappa\to K\pi$ to $\kappa\to K\eta$. Hence, by replacing these couplings we can obtain the amplitudes $\mathcal{M}^{C}_{\eta\eta}$ and $\mathcal{M}^{N}_{\eta\eta}$ for the $\eta\eta$ final states. Note that the charge conjugated loop can contribute for both the neutral and charged $\kappa$ components in this channel. The decay width of the $\eta\eta$ channel is given by
@@
$\Phi_{\eta\eta}$ is the phase space of $\eta\eta$ final states and $1/2$ is the symmetry factor.
@@
The coupling $g_{\kappa K\eta}$ can be related to $g_{\kappa^{0}K^{+}\pi^{-}}$ by the flavor symmetry. In fact, the physical $\eta$ state is the mixing of $\eta_{8}$ and $\eta_{1}$ with $\eta=\cos\theta\eta_{8}-\sin\theta\eta_{1}$ due to the SU(3) flavor breaking. As shown in Table 1, we adopt two strategies to deal with the $g_{\kappa K\eta}$ coupling: one is that we ignore the $\eta_{1}$ component in $\eta$ by taking $\eta=\cos\theta\eta_{8}$, and restrict to the symmetry inside the octet by using $P_{8}$ as the pseudoscalar matrix. In this case, we denote this coupling as $g_{\kappa K\eta}$; the other case is that we take $\eta=\cos\theta\eta_{8}-\sin\theta\eta_{1}$ and relate the couplings associated with these two part as well as the $g_{\kappa K\pi}$ by the U(3) flavor symmetry, where the pseudoscalar matrix is taken as $U$. In this case, this coupling is denoted as $g^{\prime}_{\kappa K\eta}$. The explicit forms of $P_{8}$ and $U$ can be found in Appendix. C. One must account for the fact that the $U(3)$ flavor symmetry is not good enough in the pseudoscalar sector due to the U(1) anomaly.
@@
Same as the case in $\pi\pi$ channel, there is no charge conjugated loop that can contribute. Considering the relation between the couplings $g_{{\kappa^{0}}K^{+}\pi^{-}}$ and $g_{{\kappa^{+}}K^{+}\pi^{0}}$, we find that $\mathcal{M}^{C}_{K^{+}K^{-}}=(1/2)\mathcal{M}^{N}_{K^{+}K^{-}}$. The amplitude of $f_{0}(1370)$ decay to $K^{+}K^{-}$ channel by exchanging $\eta$ is written as
@@
where the $\Phi_{K\bar{K}}$ indicate the $K\bar{K}$ phase space. The factor $2$ originates from the inclusion of $K^{0}\bar{K}^{0}$ final state.
For the $4\pi$ decay channel, there are three different specific final states which are $2\pi^{+}2\pi^{-}$, $\pi^{+}\pi^{-}2\pi^{0}$ and $4\pi^{0}$ respectively. As claimed by the analyses in experiments [20, 23, 24], the $4\pi$ final states are mainly from the intermediate channels of $\sigma\sigma$ and $\rho\rho$. This mechanism is shown in Fig. 1(c). Specifically, for the decay of an isoscalar state $f_{0}$, the $2\pi^{+}2\pi^{-}$ channel is only from the $\rho^{0}\rho^{0}$ and $\sigma\sigma$ intermediate channels; the $\pi^{+}\pi^{-}2\pi^{0}$ channel is only from the $\rho^{+}\rho^{-}$ and $\sigma\sigma$ intermediate channels; and the $4\pi^{0}$ channel is only from the $\sigma\sigma$ intermediate channel. Because these final states include identical particles, we must account for additional contributions arising from identical particles exchange.
Firstly we focus on the decay width of the $2\pi^{+}2\pi^{-}$ channel, then the decay widths of the other two channels can be related to $\Gamma_{2\pi^{+}2\pi^{-}}$ by considering the relations among the couplings in the isospin limit. As shown in Fig. 1(c), we consider the decay mechanism of $\kappa\bar{\kappa}$ rescattering to $\rho\rho$ or $\sigma\sigma$ first and subsequently decaying to $4\pi$. The amplitude of the process for $\kappa\bar{\kappa}$ rescattering to $2\sigma\to 2\pi^{+}2\pi^{-}$ by exchanging scalar $\kappa$, denoted as $\mathcal{M}^{S}_{\sigma}$, has the form of:
@@
momenta of the final $\pi$ are denoted as $\pi^{+}({p_{a}})\pi^{-}(p_{b})\pi^{+}(p_{c})\pi^{-}(p_{d})$, and $p_{ab}=p_{a}+p_{b}$, $p_{cd}=p_{c}+p_{d}$ are respectively the momenta of the two intermediate $\sigma$. The $ex.$ term indicates the contribution from exchanging the identical particles, which is the same as the first term but only change the momenta of the two $\sigma$ to $p_{ad}=p_{a}+p_{d}$ and $p_{bc}=p_{b}+p_{c}$ respectively. Note that, for such a decay mechanism, there are two charge conjugated loops for both the charged and neutral $\kappa\bar{\kappa}$ rescattering. By analyzing the relations of the couplings involved, we find that the amplitudes of these four diagrams are identical, resulting in a multiplicative factor of $4$.
@@
The amplitude of the process for the $\kappa\bar{\kappa}$ rescattering to $2\rho^{0}\to 2\pi^{+}2\pi^{-}$ by exchanging $\kappa$, as shown in Fig. 1(c), has the form of
@@
The amplitude of the process for the $\kappa\bar{\kappa}$ rescattering to $2\rho^{0}\to 2\pi^{+}2\pi^{-}$ by exchanging $K^{*}$, as shown in Fig. 1(c), has the form of
@@
here $G=g_{F\kappa\bar{\kappa}^{0}}\cdot g^{2}_{\rho^{0}\kappa^{0}\bar{K}^{*0}}\cdot g^{2}_{\rho^{0}\pi^{+}\pi^{-}}$. The total amplitude of decaying to $2\pi^{+}2\pi^{-}$ is the sum of all the amplitudes above, i.e.,
@@
The constant $1/4$ is the symmetry factor on account of the two kinds of identical particles: two $\pi^{+}$ and two $\pi^{-}$.
Then, we can discuss the decay amplitudes of $\pi^{+}\pi^{-}2\pi^{0}$ and $4\pi^{0}$ states based on the amplitude of the $2\pi^{+}2\pi^{-}$ state. For the $\rho^{+}\rho^{-}\to\pi^{+}(p_{a})\pi^{0}(p_{b})$-$\pi^{-}(p_{c})\pi^{0}(p_{d})$ process, we find the following similarities and differences compared to the $\rho^{0}\rho^{0}\to\pi^{+}(p_{a})\pi^{-}(p_{b})$-$\pi^{+}(p_{c})\pi^{-}(p_{d})$:
  • There is no charge conjugated loop for both the charged and neutral $\kappa\bar{\kappa}$ rescattering.
@@
Considering the above items we find that the amplitudes adapted to the $\rho^{+}\rho^{-}\to\pi^{+}\pi^{-}2\pi^{0}$ process are just $\mathcal{M}^{S}_{\rho}$ and $\mathcal{M}^{V}_{\rho}$.
For the $\sigma\sigma\to\pi^{+}(p_{a})\pi^{-}(p_{c})$-$\pi^{0}(p_{b})\pi^{0}(p_{d})$ process, we find the following discrepancies compared to the $\sigma\sigma\to\pi^{+}(p_{a})\pi^{-}(p_{b})$-$\pi^{+}(p_{c})\pi^{-}(p_{d})$:
  • The exchange of the two identical $\pi^{0}$ from the same $\sigma$ is topologically identical.
  • The relation of couplings: $g_{\sigma\pi^{+}\pi^{-}}=2g_{\sigma\pi^{0}\pi^{0}}$.
  • There are two ways to contract the two $\pi^{0}$ fields and the two final $\pi^{0}$ states at the $\sigma\pi^{0}\pi^{0}$ vertex.
Therefore, the amplitudes $\mathcal{M}^{S}_{\sigma}$ and $\mathcal{M}^{V}_{\sigma}$ adapted to the $\pi^{+}\pi^{-}2\pi^{0}$ state are obtained by changing the momenta of the two $\sigma$ to $p_{ac}=p_{a}+p_{c}$ and $p_{bd}=p_{b}+p_{d}$ respectively and eliminating the exchange terms.